OpenStudy (anonymous):

A particle moves under a hypothetical force so that its velocity vector is v= . Find its radial acceleration...Please help...

OpenStudy (anonymous):

whata hard bout that?

OpenStudy (anonymous):

acceleration is the derivative of velocity with respect to time

OpenStudy (anonymous):

assuming k is a constant

OpenStudy (anonymous):

\[a = < k \frac{d}{dt} ( tcos(t)) , k \frac{d}{dt} ( tsin(t) ) > \]

OpenStudy (anonymous):

k can be factored out of the vector, and it will just provide a scalar multiple

OpenStudy (anonymous):

\[a = k < ( \cos(t) -tsin(t)) , (\sin(t) + tcos(t)) > \]

OpenStudy (anonymous):

by product rules

OpenStudy (anonymous):

wow you are such a genius. I admire you all

OpenStudy (anonymous):

?

OpenStudy (anonymous):

bt i need the radial accn, not the total accn....

OpenStudy (anonymous):

at time t=0

OpenStudy (anonymous):

yeh i dnt have clue on that :|

OpenStudy (anonymous):

radial is apparently equal to centripetal= v^2 / r

OpenStudy (anonymous):

but i dont have a clue

OpenStudy (anonymous):

hey i got it neway ..thnx

OpenStudy (nikvist):

\[\vec{a}_n=<-kt\sin{t},kt\cos{t}>\]